OpenStudy (anonymous):

solve. sqrt(x+5)-1=sqrt(x)

OpenStudy (anonymous):

\[\sqrt{x+5}-1=\sqrt{x}\]

OpenStudy (anonymous):

\[\sqrt{x + 5} - 1 = \sqrt{x}\]

OpenStudy (anonymous):

Take 1 to the other side and square on both sides.

OpenStudy (anonymous):

ok so i got: \[x+5=x+2\sqrt{x}+1\]

OpenStudy (anonymous):